OpenStudy (anonymous):

Medal&Fan Joel has 2/3 of a large candy bar that he plans to divide into 4 pieces for himself and his three friends. What fraction of the original large candy bar will each person get? Select one: a. 3/8 b. 8/3 c. 1/6 d. 1/12

OpenStudy (hoblos):

\[\frac{ 2 }{ 3 }\div4 = \frac{ 2 }{ 3\times4 } = \frac{ 2 }{ 12 } = \frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

its c 1/6

OpenStudy (anonymous):

thanks guys!

OpenStudy (hoblos):

any time :)